### v1

Abstract
Large Language Models (LLMs) have rapidly advanced, with Gemini-3-Pro setting a new performance milestone. In this work, we explore collective intelligence as an alternative to monolithic scaling, and demonstrate that open-source LLMs’ collaboration can surpass Gemini-3-Pro. We first revisit LLM routing and aggregation at scale and identify three key bottlenecks: (1) current train-free routers are limited by a query-based paradigm focusing solely on textual similarity; (2) recent aggregation methods remain largely static, failing to select appropriate aggregators for different tasks; (3) the complementarity of routing and aggregation remains underutilized. To address these problems, we introduce JiSi, a novel framework designed to release the full potential of LLMs’ collaboration through three innovations: (1) Query-Response Mixed Routing capturing both semantic information and problem difficulty; (2) Support-Set-based Aggregator Selection jointly evaluating the aggregation and domain capacity of aggregators; (3) Adaptive Routing-Aggregation Switch dynamically leveraging the advantages of routing and aggregation. Comprehensive experiments on nine benchmarks demonstrate that JiSi can surpass Gemini-3-Pro with only 47% costs by orchestrating ten open-source LLMs, while outperforming mainstream baselines. It suggests that collective intelligence represents a novel path towards Artificial General Intelligence (AGI).
Machine Learning, ICML
Figure 1: The overall leaderboard of different LLMs and the proposed JiSi. By orchestrating open-source LLMs collectively, JiSi surpasses all monolithic LLMs, including leading closed-source LLMs such as Gemini-3-Pro, suggesting that collective intelligence represents a novel path towards AGI.
1 Introduction
The landscape of Artificial Intelligence (AI) has been fundamentally reshaped by Transformer-based Large Language Models (LLMs) (vaswani2017attention; brown2020language). By continuously scaling up model parameters and training data, LLMs have achieved unprecedented capabilities (kaplan2020scaling; comanici2025gemini25pushingfrontier; claude-sonnet-4; gpt-5; gpt-5.2; gemini-3-pro). Most notably, Gemini-3-Pro (gemini-3-pro) has established a significant performance milestone across different fields, including knowledge and reasoning, coding and engineering, general chat and factuality. The overwhelming success of Gemini-3-Pro further reinforces the prevailing paradigm of training increasingly powerful monolithic models. However, this naturally raises a fundamental question: is the indefinite scaling of a monolithic “super model” the only path towards Artificial General Intelligence (AGI)?
(a) Comparison with open-source LLMs.
(b) Comparison with closed-source LLMs.
Figure 2: Comprehensive performance comparison between the proposed JiSi and other LLMs. (a) and (b) show the fine-grained capability comparison with other open-source and closed-source LLMs on different datasets, respectively. JiSi achieves superior performance across multiple benchmarks.
In this work, we instead explore collective intelligence as an alternative path, demonstrating that the collaboration of individually weaker open-source LLMs can surpass leading closed-source models, including Gemini-3-Pro. Existing approaches to multi-LLM collaboration can be broadly divided into two fundamental categories. The first category is routing-based methods. Representative approaches include RouterDC (chen2024routerdc), EmbedLLM (zhuang2024embedllmlearningcompactrepresentations), GraphRouter (feng2025graphroutergraphbasedrouterllm), Avengers (zhang2025avengerssimplerecipeuniting), and retrieval-based routers (tang2025opensourcellmscollaborationbeats). These methods typically rely on query embeddings based similarity matching, or lightweight performance predictors to select the most suitable model for each input. The second category is aggregation-based methods. Typical techniques include mixture-of-agents (wang2024mixtureofagentsenhanceslargelanguage), self-consistency (chen2023universalselfconsistencylargelanguage), and majority voting (chen2024llmcallsneedscaling). These methods seek to surpass the limitations of individual models by synthesizing complementary information or consolidating consensus from multiple LLMs.
To achieve our goal, we first revisit routing and aggregation at scale, under both giant LLMs and challenging benchmarks. We find that: (1) current routing methods (shnitzer2023large; chen2024routerdc; zhang2025avengerssimplerecipeuniting; zhuang2024embedllmlearningcompactrepresentations) are predominantly training-free, typically projecting queries into a latent space for clustering or matching based model selection. However, pre-trained embedding models only capture text similarity and fail to perceive deep semantics or task difficulty, leading to suboptimal or mismatched selection. (2) Recent aggregation techniques (wang2024mixtureofagentsenhanceslargelanguage; li2025rethinking; li2025smoa) mainly exploit complementary information but are largely static, making it difficult to adaptively select aggregators with strong aggregation and domain-specific capabilities, which inevitably suffers from bottlenecks at critical responses generation. (3) the complementarity between routing and aggregation is often overlooked: routing offers stability but is constrained by single model, aggregation can exceed individual model limits yet is vulnerable to low-quality outputs and noise (li2025smoa; wolf2025your; li2025rethinking), while straightforward combination approaches struggle to effectively leverage their respective strengths.
To break through above bottlenecks, we propose a novel framework entitled JiSi [^1] JiSi is named after an idiom that conveys the idea that collective intelligence emerges from aggregating diverse perspectives. The project will be updated at https://github.com/magent4aci/openJiSi., featuring a minimalist but surprisingly effective designs. Specifically, based on a pre-built embedding bank, it contains three core techniques: 1) Query-Response Mixed Routing: To capture deep semantics and task difficulty beyond shallow text embeddings, we utilize LLM-generated response embeddings to refine the routing process, causing a more precise match between input queries and LLMs. 2) Support-Set-based Aggregator Selection: To strike a balance between domain-specific and general aggregation capabilities, we leverage prior scores derived from a larger-scale embedding support set to select suitable aggregators dynamically. 3) Adaptive Routing-Aggregation Switch: We introduce an adaptive switch mechanism that selectively aggregates excellent expert responses or switches to a pure routing strategy according to their refined prior scores. This strategy suppresses noise and prevents the propagation of inferior information to the aggregator.
We conduct comprehensive experimental comparisons and analyses to assess the effectiveness of the proposed method JiSi. Specifically, we select ten open-source giant LLMs as candidate models and evaluate JiSi across nine challenging benchmarks. As shown in Fig. 1, the proposed JiSi surpasses all current state-of-the-art closed-source proprietary models. Compared to Gemini-3-Pro, it achieves an average performance gain of +1.15 and saves 53.23% cost. As shown in Fig. 2, JiSi can take specific advantages of candidate LLMs and break through their upper bound. In addition, JiSi remarkably outperforms all mainstream baselines. Furthermore, we verify that JiSi, as a concise training-free framework, exhibits consistent performance improvement with the introduction of new LLMs, which highlights its strong scalability. These results suggest that collective intelligence is a potential new path towards AGI.
2 Related works
Collective intelligence has emerged as a promising paradigm for pushing the intelligence upper bound. By orchestrating collaboration mechanisms across multiple models, collective intelligence aims to substantially enhance performance on complex tasks. This paper investigates collective intelligence in multi-LLM collaboration, which can be broadly divided into routing-based and aggregation-based methods.
Router-based Methods. These approaches exploit the domain-specific strengths of heterogeneous models by routing queries based on task requirements. For instance, ZOOTER (lu2024routing) proposed to distill signals from off-the-shelf reward models to train a router capable of precisely distributing queries to LLMs with corresponding expertise. RouterDC (chen2024routerdc) enhances routing accuracy through dual contrastive learning. GraphRouter (feng2025graphroutergraphbasedrouterllm) constructs a heterogeneous task-query-LLM graph, formulating model selection as a dynamic edge prediction problem via GNN. Similarly, MODEL-SAT (zhang2025capability) encodes candidate model performance on core tasks into capability representations, utilizing a lightweight LLM to predict the optimal candidate. Avengers (zhang2025avengerssimplerecipeuniting) constructs a training-free router by clustering the question embeddings as the proxies of different question categories, and utilizing the performance of the clusters as a model profile to estimate the LLM candidates.
Aggregation-based Methods. These methods focus on integration, synthesizing independent outputs from multiple LLMs to generate a superior response. Agent-Forest (li2024more) employs a majority voting mechanism to select the most frequent consensus among responses. The Mixture of Agents (MoA) framework (wang2024mixtureofagentsenhanceslargelanguage) introduces a layered architecture that leverages the collective strengths of LLMs through iterative collaboration, significantly improving generation quality. Building on this, SMoA (li2025smoa) incorporates response selection and early stopping mechanisms to sparsify information flows among agents, balancing performance with efficiency. Self-MoA (li2025rethinking) explores aggregating responses repeatedly sampled from a single high-performance agent.
[…]
Figure 3: JiSi rethinks the existing routing and aggregation methods, and reshapes them from three aspects: 1) Routing: from query-based to query-response-mixed; 2) Aggregation: from fixed aggregator to support-set-based aggregator selection; 3) Combination: from static routing and aggregation to adaptive routing-aggregation switch. For simplicity, the “Agg” means Aggregator or Aggregation.
3 JiSi Framework
The JiSi framework stems from the rethinking w.r.t. existing routers and aggregation methods. In Fig. 3, we intuitively display the core improvement of JiSi compared with the previous methods in three aspects.
Rethinking Routers. Current LLM routers typically adopt a query-based scheme to select an LLM. To evaluate the connection between the given query and candidate LLMs, they directly extract the query representation via pre-trained embedding models. The query representation is used for clustering to build LLM profiles (zhang2025avengerssimplerecipeuniting; chen2024routerdc) or training LLM embedding (zhuang2024embedllmlearningcompactrepresentations). However, because the pre-trained embedding models are trained for general-purpose text retrieval, the extracted embedding focuses on the query text similarity while ignoring the deep essential information beyond text overlap, such as deep semantics and question difficulty. For example, “Prove every large even number is a sum of two primes” and “Prove every large even number is a sum of two odds” almost have the same words, except for the last word. However, the first question is far more difficult than the second one. Based on query embeddings, they are hardly distinguished. It results in suboptimal query features or corresponding query support sets, which leads to routing these two questions to the same LLM incorrectly.
To address this issue, JiSi leverages the potent semantic extraction capabilities of LLMs to obtain auxiliary information from both generated responses and token costs. Specifically, the response to a query provides latent semantic information beyond surface-level text similarity. For instance, an LLM may generate significantly different responses to the two aforementioned questions, enabling the embedding model to effectively distinguish them. Moreover, based on an intuitive heuristic that a more difficult query typically correlates with more token costs for the same LLM, JiSi incorporates token costs to achieve a more precise calibration of query difficulty. Notably, query text similarity remains indispensable as it captures essential coarse-grained semantic features. Thus, by a weighted integration of the query, response, and token costs, JiSi achieves more accurate model routing.
Rethinking Aggregation. Existing aggregation methods (li2025smoa; tang2025opensourcellmscollaborationbeats) are heuristic and statically select an aggregator based on the overall performance, which focuses on the comprehensive ability of aggregators. It may not be suitable for a specific task. However, while an LLM’s aggregation proficiency is inherently correlated with its comprehensive capability, a domain-specific expert does not necessarily possess superior aggregation skills. Consequently, the ideal aggregator must strike a balance between domain-specific and comprehensive ability. To identify the optimal task-specific aggregator, JiSi evaluates candidates by considering both their comprehensive performance across broader queries and their specialized expertise within the specific knowledge domain of the given query. It is worthnoting that since Fig. 3 aims to give an intuitive visualization of the proposed method, the practical deployment is different. JiSi utilizes the coarse-grained characteristic of query text similarity. It constructs a larger query support set in a query bank via query-based similarity that is rich in task information but contains diverse categories of queries, which guarantees the balanced estimation of domain-specific expertise and comprehensive ability. In this manner, JiSi can dynamically identify a domain-specific and synthesis-skilled LLM for a given new query.
Rethinking Combination. Existing methods (chen2025symbolic; tang2025opensourcellmscollaborationbeats) generally adopt an inflexible combination of routing and aggregation, typically following “routing many and aggregating all” workflow that treats aggregation as a post-processing step of routing. However, the complementarity and dynamics of routing and aggregation have long been overlooked. Routers may be advantageous for some simple and direct tasks, whereas aggregation methods can be useful for other complex tasks. Thus, always applying aggregation is not efficiency with a large range of query complexity. Moreover, while routers are more stable because it only selects high-quality LLMs, it hardly breaks through the upper bound of single-model performance; conversely, aggregation can exceed the capability of individual models but suffers from inferior or noisy inputs. To synergize the advantages of routing and aggregation, we propose an adaptive routing-aggregation switch mechanism for output strategies based on prior scores. Specifically, JiSi utilizes a pre-built embedding bank during the routing phase to obtain a precise prior score for each candidate response. It prompts the pruning of inferior or noisy responses and allows the system to adaptively switch between aggregation modes of varying input numbers. Notably, if only a single response exceeds the high-score threshold, the strategy bypasses the aggregation process and applies direct routing, significantly enhancing inference efficiency.
3.1 Framework Overview
In summary, JiSi adopts a minimalist “route-and-aggregate” topology: based on a pre-built embedding bank, it conducts a unified LLM routing and responses aggregation. For embedding bank construction, an extensible set of queries is sampled from heterogeneous sources. Responses from candidate LLMs are collected for these questions. Afterwards, all queries and responses are fed into a pre-trained embedding model to obtain latent representations, which form the embedding bank. Based on the embedding bank, the processing of a new query comprises three essential components. 1) Support-Set-based Aggregator Selection: A support query set is selected for aggregator dispatching and then filtered to improve the assignment of candidate aggregatee models. 2) Query-Response Mixed Routing: For each model selected from the filtered support set, a composite score is derived from query, response, and reasoning-length similarities, serving as a prior evaluation of the model’s expertise capacity for the given query. 3) Adaptive Routing-Aggregation Switch: the practical aggregatees are determined based on their scores and response length to filter inferior responses and noise. If the number of aggregatees exceeds 1, aggregation is performed to obtain the final response; otherwise, the final response is retrieved from the single aggregatee, i.e., it falls back to top-1 routing. An overview of the framework is illustrated in Fig. 4.
Figure 4: Overview of the proposed JiSi. For a new question query, a support set is first extracted from the embedding bank based on question similarity. Consequently, three essential techniques, Support-Set-based Aggregator Selection, Query-Response Mixed Routing, and Adaptive Routing-Aggregation Switch, are combined to select and aggregate LLMs dynamically.
3.2 Embedding Bank
The embedding bank is constructed as an LLM profile storage, facilitating architecture-agnostic evaluation. Intuitively, the embedding bank is built on an extensive collection of queries from diverse sources. The original queries, LLM responses, and correctness information are subsequently projected into continuous embedding spaces to support similarity analysis.
Specifically, let $\left[U\right]=\left[1,U\right]\cap\mathbb{N}$. Assuming that $N$ questions $\mathcal{Q}=\left\{\left.x_{i}\in\mathcal{S}\right|i\in\left[N\right]\right\}$ each with corresponding label $y_{i}\in\mathcal{S}$, and $M$ LLMs $\mathcal{A}=\left\{\boldsymbol{A}_{1},\boldsymbol{A}_{2},\dots,\boldsymbol{A}_{M}\right\},\,\boldsymbol{A}_{i}:\mathcal{S}\to\mathcal{S}$ are predetermined. The $\mathcal{S}$ contains all literal strings. Responses of each LLM are collected over all questions, obtaining $\mathcal{R}=\left\{\left.\hat{y}_{i,j}=\boldsymbol{A}_{j}\left(x_{i}\right)\right|i\in\left[N\right],j\in\left[M\right]\right\}$ with reasoning costs (completion tokens) $c_{i,j}\in\mathbb{R}_{+}$. A pre-trained embedding model $\boldsymbol{A}_{emb}:\mathcal{S}\to\mathbb{R}^{d}$ is introduced to map questions and responses into contiguous latent space, forming the embedding bank $\mathcal{E}$:
$\mathcal{E}=\mathcal{E}^{Q}\cup\mathcal{E}^{R},$ (1)
$\mathcal{E}^{Q}=\left\{\boldsymbol{e}_{i}=\left.\boldsymbol{f}_{1}\left(x_{i}\right)\right|i\in\left[N\right]\right\},$
$\mathcal{E}^{R}=\left\{\boldsymbol{r}_{i,j}=\left.\boldsymbol{f}_{1}\left(\hat{y}_{i,j}\right)\right|i\in\left[N\right],j\in\left[M\right]\right\},$
$\boldsymbol{f}_{1}(a)=\frac{\boldsymbol{A}_{\mathrm{emb}}(a)}{\|\boldsymbol{A}_{\mathrm{emb}}(a)\|_{2}}\in\mathbb{R}^{d},a\in\mathcal{S},$
which can be splitted into question embedding set $\mathcal{E}^{Q}$ and response embedding set $\mathcal{E}^{R}$. Moreover, for each LLM $\boldsymbol{A}_{j}$, a capability vector $\boldsymbol{v}_{j}\in\left\{0,1\right\}^{N}$ is generated to measure response correctness over questions, i.e., $\boldsymbol{v}_{j}=\left[\boldsymbol{1}_{\left\{y_{1}\right\}}\left(\hat{y}_{1,j}\right),\dots,\boldsymbol{1}_{\left\{y_{N}\right\}}\left(\hat{y}_{N,j}\right)\right]^{T}$, where $\boldsymbol{1}_{A}\left(\cdot\right)$ is a general indicator function, meaning computing the correctness.
3.3 Inference
Support-Set-based Aggregator Selection. Provided a query denoted as $x$. Its embedding can be extracted as $\boldsymbol{e}=\boldsymbol{f}_{1}\left(x\right)$. Similarities between $\boldsymbol{e}$ and all embeddings in $\mathcal{E}^{Q}$ are then calculated via cosine distance (note that the embeddings are already normalized): $\boldsymbol{s}=\left[\boldsymbol{e}_{1}^{T}\boldsymbol{e},\dots,\boldsymbol{e}_{N}^{T}\boldsymbol{e}\right]^{T}\in\left[0,1\right]^{N}$, where $\boldsymbol{e}_{i}\in\mathcal{E}^{Q},\,i\in\left[N\right]$. To determine the support question number $N^{sup}$, a predefined base number $N^{base}$ is used. The indices included in the support set are then derived as $\mathcal{I}=\left\{i_{1},\dots,i_{N^{sup}}\right\}=\mathcal{I}\left\{\left.i\in\left[N\right]\right|s_{i}\geq\gamma\left[\boldsymbol{f}_{2}\left(N^{base}\right)\right]\left(\boldsymbol{s}\right)\right\}$, where $\gamma\in\left[0,1\right]$ is predefined and
$\boldsymbol{f}_{2}:$ $\left[U\right]\to\mathcal{F},\,\mathcal{F}=\left\{\left.\boldsymbol{f}\right|\boldsymbol{f}:\mathbb{R}^{U}\to\mathbb{R}\right\},$ (2)
[…]
$\operatorname{s.t.}$ $\left|\left\{\left.i\in\left[U\right]\right|a_{i}\geq\hat{a}\right\}\right|=\alpha,$
[…]
Consequently, $N^{sup}$ can be obtained via $\left|\mathcal{I}\right|$. Applying $\mathcal{I}$ to both $\boldsymbol{s}$ and the concatenated $\boldsymbol{v}=\left[\boldsymbol{v}_{1}^{T},\dots,\boldsymbol{v}_{M}^{T}\right]^{T}\in\left\{0,1\right\}^{M\times{N}}$ results in $\boldsymbol{s}^{sup}=\boldsymbol{s}\left[\mathcal{I}\right]\in\left[0,1\right]^{N^{sup}}$ and $\boldsymbol{v}^{sup}=\boldsymbol{v}\left[:,\mathcal{I}\right]\in\left\{0,1\right\}^{M\times{N^{sup}}}$. The vector slicing operator is denoted as $\left[\cdot\right]$. The coarse-grained model score $\boldsymbol{g}\in\mathbb{R}_{+}^{M}$ is computed via $\boldsymbol{v}^{sup}\boldsymbol{s}^{sup}$. The aggregator $\boldsymbol{A}_{agg}=\boldsymbol{A}_{k}$ is selected, where $g_{k}=\left[\boldsymbol{f}_{2}\left(1\right)\right]\left(\boldsymbol{g}\right)$. The support set selected provides a broader knowledge base than the subsequent filtered set. The aggregator is anchored in this larger set to ensure a better understanding of divergent domains, thereby enabling more skilled aggregation.
[…]
$\boldsymbol{s}^{flt}$ $=\left(\epsilon{K}\boldsymbol{s}+\sigma\boldsymbol{s}^{res}+\delta\boldsymbol{s}^{cost}\right)\left[\mathcal{I}\right],$
$s_{i}^{res}$ $=\sum\nolimits_{j=1}^{K}\boldsymbol{r}_{i,j}^{T}\boldsymbol{r}_{j},\,\boldsymbol{r}_{i,j}\in\mathcal{E}^{R},\,i\in\left[N\right],$
$s_{i}^{cost}$ $=\sum\nolimits_{j=1}^{K}\left(1-\left(\left|c_{i,j}^{2}-c_{j}^{2}\right|/c_{m,j}\right)\right),\,i\in\left[N\right],$
$c_{m,j}$ $=\left[\boldsymbol{f}_{2}\left(1\right)\right]\left(\left[\left|c_{i_{1},j}^{2}-c_{j}^{2}\right|,\dots,\left|c_{i_{N^{sup}},j}^{2}-c_{j}^{2}\right|\right]\right),$
$\operatorname{s.t.}\ \left[\beta,\epsilon,\sigma,\delta\right]\in\left[0,1\right]^{4},\,\epsilon+\sigma+\delta=1.$
The hyperparameters $\beta,\,\epsilon,\,\sigma,\,\delta$ are predefined. The filtered indices account for question similarity, model response similarity, and reasoning cost similarity, leading to a more precise support set for better aggregatee selection. With respect to $\mathcal{I}^{flt}$, a fine-grained model score is computed as $\boldsymbol{g}^{f}=\boldsymbol{v}^{sup}\left[:,\mathcal{I}^{flt}\right]\boldsymbol{s}^{flt}\left[\mathcal{I}^{flt}\right]\in\mathbb{R}_{+}^{N}$. Compared with the query-only support set, the filtered set accounts for both queries and LLM responses, helping to filter queries that are semantically similar but differ in difficulty, which is captured by the deep semantics and length of LLM responses.
Adaptive Routing-Aggregation Switch. To conduct adaptive aggregation, a threshold $t\in\left[0,1\right]$ is predetermined. The final aggregatees $\mathcal{A}^{fm}$ are formulated as:
$\mathcal{A}^{fm}$ $=\biggl\{\left.\boldsymbol{A}_{i}\in\mathcal{A}\right|i\in\left[M\right],$ (4)
[…]
Let $N^{fm}$ being $\left|\mathcal{A}^{fm}\right|$ and the elements in $\mathcal{A}^{fm}$ being $\boldsymbol{A}_{i}^{fm}$, the final response $\hat{y}$ is gathered via:
[…]
The $\boldsymbol{f}_{3}:2^{\mathcal{S}}\to\mathcal{S}$ concates any number of strings to a single string. When the number of aggregatees equals 1, aggregation falls back to top-1 fine-grained model score routing. Otherwise, the aggregation dynamically selects LLMs (at most K) to provide context. Note that, to obtain $\bar{x}$, the responses $\hat{y}_{k},\,k\in\left[K\right]$ can be mostly reused due to the overlapping between $\mathcal{A}^{cm}$ and $\mathcal{A}^{fm}$. The adaptive switch fully leverages the advantages of routing and aggregation by filtering noisy LLM responses using a threshold, thereby guaranteeing high-quality aggregation when triggered; otherwise, it provides a more efficient direct-routing path that is less distracting.
Model AIME Arena-Hard GPQA HLE LiveCodeBench LiveMathBench MMLU-Pro SimpleQA $\text{SWE-bench}^{\dagger}$ Avg
Open-source LLMs
DeepSeek-R1-0528 (deepseekai2025deepseekr1incentivizingreasoningcapability) 72.22 64.89 78.33 16.67 76.03 72.97 84.67 28.66 25.33 57.75
DeepSeek-V3-0324 (deepseekai2024deepseekv3technicalreport) 38.89 59.56 68.33 3.70 61.51 59.46 78.44 26.43 24.00 46.70
DeepSeek-V3.1-Terminus (deepseekai2024deepseekv3technicalreport) 55.56 64.67 78.33 8.64 64.67 67.57 84.56 25.12 26.00 52.79
GLM-4.6 (glm-4.6) 88.89 69.56 80.00 14.20 58.99 64.86 80.89 25.89 22.67 56.22
Intern-S1 (bai2025interns1scientificmultimodalfoundation) 38.89 68.00 70.00 9.72 46.69 59.46 83.00 14.33 8.00 44.23
Kimi-K2-0905 (kimiteam2025kimik2openagentic) 72.22 72.22 71.67 5.09 62.15 75.68 80.78 30.66 24.00 54.94
DeepSeek-V3.2-Thinking (deepseekai2025deepseekv32pushingfrontieropen) 88.89 62.44 88.33 24.69 83.91 78.38 87.33 27.81 24.67 62.94
DeepSeek-V3.2-Speciale (deepseekai2025deepseekv32pushingfrontieropen) 94.44 55.33 83.33 27.16 86.75 75.68 87.44 39.52 40.67 65.59
Qwen3-235B-A22B-2507 (yang2025qwen3technicalreport) 77.78 75.33 55.00 9.41 58.36 72.97 83.78 54.01 16.67 55.92
Qwen3-235B-A22B-Thinking-2507 (yang2025qwen3technicalreport) 72.22 77.78 80.00 7.56 75.71 48.65 80.56 49.31 20.00 56.87
Closed-source LLMs
Claude-Sonnet-4 (claude-sonnet-4) 41.11 55.47 71.33 4.60 56.85 62.16 83.58 15.58 35.33 47.33
Claude-Sonnet-4.5 (claude-sonnet-4.5) 27.78 64.00 71.67 7.56 60.57 59.46 86.33 16.18 34.00 47.51
Gemini-2.5-Pro (comanici2025gemini25pushingfrontier) 82.22 77.64 84.00 20.62 77.54 45.95 86.69 54.33 34.53 62.61
Grok-4 (grok-4) 88.89 56.89 88.33 24.42 81.03 75.68 86.56 48.38 27.33 64.17
GPT-5 (gpt-5) 83.33 67.11 88.33 25.77 84.54 78.38 87.22 48.00 16.00 64.30
GPT-5.2-Thinking (gpt-5.2) (High Effort) 83.33 85.78 93.33 29.94 90.50 78.38 86.67 35.21 12.67 66.20
Gemini-3-Pro (gemini-3-pro) 94.44 74.55 91.67 33.02 89.59 78.38 89.33 70.03 18.00 71.00
Router Methods
RouterDC (chen2024routerdc) 88.89 76.00 73.33 21.30 83.60 78.38 84.33 53.31 24.00 64.79
EmbedLLM (zhuang2024embedllmlearningcompactrepresentations) 94.44 57.78 90.00 27.31 84.86 75.68 87.44 53.70 42.67 68.21
GraphRouter (feng2025graphroutergraphbasedrouterllm) (Performance-only) 88.89 82.67 90.00 25.93 86.12 78.38 87.33 49.08 27.33 68.41
Avengers (zhang2025avengerssimplerecipeuniting) 94.44 70.44 83.33 27.16 85.49 75.68 87.33 54.08 40.67 68.74
Retrieval-based Router (tang2025opensourcellmscollaborationbeats) 94.44 72.67 86.67 26.85 85.17 75.68 86.78 53.78 40.00 69.12
Our Query-Response Mixed Router 94.44 76.89 83.33 26.70 85.17 75.68 87.00 54.62 43.33 69.68
Multi-agent Methods
Best (open-source) LLM 94.44 77.78 88.33 27.16 86.75 78.38 87.44 54.01 40.67 70.55
MOA (wang2024mixtureofagentsenhanceslargelanguage) (3 Best LLMs) 94.44 80.00 78.33 28.70 88.33 75.68 86.44 42.99 33.33 67.58
Self Consistency (chen2023universalselfconsistencylargelanguage) (Best LLM with 4 Rollouts) 94.44 83.70 85.00 27.62 88.96 75.68 87.56 54.31 38.67 70.66
Majority Voting (chen2024llmcallsneedscaling) (10 LLMs) 83.33 74.67 85.00 13.58 72.56 72.97 87.44 36.98 28.67 61.69
Majority Voting (chen2024llmcallsneedscaling) (3 Best LLMs) 88.89 72.44 86.67 25.93 85.80 75.68 86.78 30.20 29.33 64.64
Our JiSi (w/o Adaptive Aggregation) 94.44 86.44 85.00 30.09 89.27 78.38 87.44 51.46 37.33 71.09
Our JiSi 94.44 88.44 86.67 27.62 89.27 81.08 86.78 53.7 41.33 72.15
Table 1: Main results over diverse benchmarks, LLMs, and methods. The proposed JiSi exhibits superior performance, ranking the best among all open-source and closed-source LLMs, routing methods, and multi-agent methods. $\text{SWE-bench}^{\dagger}$: following the evaluation in (openrouterbench), we pack all context in a single query and drive the LLM or system to solve the issue only via a single response
4 Experiment
In this section, a comprehensive evaluation of the proposed framework is presented, involving nine benchmarks, fifteen LLMs, five router baselines, and five multi-agent baselines.
4.1 Settings
Data Preparation. To verify the effectiveness of the proposed framework across benchmarks and different LLM compositions, experiments rely heavily on OpenRouterBench (openrouterbench) (until commit b42d10c) to form both the embedding bank and the test set. For each dataset, the data are split into a training set and a test set in a 0.7:0.3 ratio. In more detail, all data are first added to a list in ascending order by the recorded index, and then shuffled using Numpy (harris2020array) with seed 42. The training set comprises the first 0.7 proportion of the shuffled list, and the test set comprises the remainder. The following benchmarks are involved in the experiments: AIME (aime), Arena-Hard (arenahard2024), GPQA (rein2024gpqa), HLE (phan2025humanitysexam), LiveCodeBench (jain2024livecodebench), LiveMathBench (livemathbench), MMLU-Pro (wang2024mmlu), SimpleQA (wei2024measuringshortformfactualitylarge), and SWE-bench (jimenez2024swebench).
Evaluation. The evaluation follows (openrouterbench) and uses non-interactive mode with a single-round response. Specifically, the judger for Arena-Hard is DeekSeek-V3-0324 (deepseekai2024deepseekv3technicalreport), and the judger for HLE and SimpleQA is o3-mini (o3-mini). For SWE-bench, because aggregation methods are hard to conduct multi-turn interaction and tool usage, we pack all context in a single query and drive the LLM/system to solve the issue only via a single response following (openrouterbench). For an LLM, its responses on the SWE-bench test set are packaged and sent to sb-cli to finalize the results.
Implementation Details. For each item in the training set, the original question and LLM responses are extracted and fed into gte-Qwen2-7B-instruct (li2023generaltextembeddingsmultistage) to obtain embeddings. Additionally, information about completion tokens and response correctness is used to compute reasoning costs and the capability vector. If a response is missing during embedding bank construction or test set inference, an API remote call is triggered to retrieve it from LLM providers, given temperature 0.2 and top-p 1.0. The received response is then cached back to OpenRouterBench. The router and multi-agent methods are applied to ten open-source LLMs ($M=10$), including DeepSeek-R1-0528 (deepseekai2025deepseekr1incentivizingreasoningcapability), DeepSeek-V3-0324 (deepseekai2024deepseekv3technicalreport), DeepSeek-V3.1-Terminus (deepseekai2024deepseekv3technicalreport), GLM-4.6 (glm-4.6), Intern-S1 (bai2025interns1scientificmultimodalfoundation), Kimi-K2-0905 (kimiteam2025kimik2openagentic), DeepSeek-V3.2-Thinking (deepseekai2025deepseekv32pushingfrontieropen), DeepSeek-V3.2-Speciale (deepseekai2025deepseekv32pushingfrontieropen), Qwen3-235B-A22B-2507 (yang2025qwen3technicalreport), and Qwen3-235B-A22B-Thinking-2507 (yang2025qwen3technicalreport). Besides, seven closed-source LLMs are selected for comparison: Claude-Sonnet-4 (claude-sonnet-4), Claude-Sonnet-4.5 (claude-sonnet-4.5), Grok-4 (grok-4), Gemini-2.5-Pro (comanici2025gemini25pushingfrontier), GPT-5 (gpt-5), GPT-5.2-Thinking (gpt-5.2) with high thinking effort, and Gemini-3-Pro (gemini-3-pro). Without otherwise mentioned, $K,\,N^{base},\,\beta,\,\gamma,\,\epsilon,\,\sigma,\,\delta,\,t$ are set to 3, 50, 0.5, 0.95, 0.5, 0.3, 0.2, 0.8, respectively. To avoid extremely long context, the aggregatee number is truncated to 2 if the token sum of 3 aggregatees exceeds 13,000. For computing the cost, if the official API is available, we use the official API price. Otherwise, we adopt the stable API servers with the lowest price in OpenRouter [^2] https://openrouter.ai/.
AIME Arena-Hard GPQA HLE LiveCodeBench LiveMathBench MMLU-Pro SimpleQA SWE-bench All Cost (JiSi saving ratio) Avg Performance
Grok-4 \$3.98 \$11.11 \$8.97 \$171.56 \$45.78 \$2.69 \$50.08 \$22.44 \$29.24 \$345.85 ($\downarrow$ 81.68%) 64.17
GPT-5 \$1.57 \$7.97 \$2.22 \$39.29 \$18.17 \$1.43 \$10.25 \$16.98 \$7.49 \$105.37 ($\downarrow$ 39.87%) 64.30
GPT-5.2-Thinking \$1.07 \$5.39 \$1.64 \$52.49 \$15.62 \$1.18 \$6.49 \$16.80 \$11.52 \$112.20 ($\downarrow$ 43.53%) 66.20
Gemini-3-Pro \$1.73 \$4.57 \$3.18 \$52.47 \$22.91 \$1.97 \$23.29 \$7.56 \$17.78 \$135.46 ($\downarrow$ 53.23%) 71.00
JiSi (Open-source Collective) \$0.61 \$2.95 \$1.23 \$31.01 \$9.38 \$0.59 \$7.12 \$4.85 \$5.62 \$63.36 72.15
Table 2: The cost report of the proposed open-source collective JiSi and leading proprietary LLMs. The cost is measured by U.S. dollars. Compared with proprietary LLMs, JiSi can save approximately 40% $\sim$ 80% cost with better performance.
Figure 5: The scaling curve of the proposed JiSi, where we add open-source LLMs following their chronological release sequence as a simulation of the real-world evolution of the LLM ecosystem.
Figure 6: The distribution of LLM selection using the proposed JiSi on different datasets. JiSi prefers to select the high-performance LLMs while maintaining the task-aware specialization and abundant usage diversity.
4.2 Main Results
As demonstrated in Table 1, the proposed JiSi is compared with the aforementioned closed-source and open-source LLMs, along with typical router methods and multi-agent methods, i.e., RouterDC (chen2024routerdc), EmbedLLM (zhuang2024embedllmlearningcompactrepresentations), GraphRouter (feng2025graphroutergraphbasedrouterllm), Avengers (zhang2025avengerssimplerecipeuniting), Retrieval-based Router (tang2025opensourcellmscollaborationbeats), MOA (wang2024mixtureofagentsenhanceslargelanguage), Self Consistency (chen2023universalselfconsistencylargelanguage), and Majority Voting (chen2024llmcallsneedscaling). For multi-agent baselines, priors from OpenRouterBench are applied to provide a stronger variant. For each benchmark, the MOA selects three LLMs with the highest evaluation scores over the training records on the benchmark; the Self Consistency sample four rollouts from the best LLM evaluated on training data and merges responses via majority voting; the Majority Voting is verified in two settings. The first is vanilla, encompassing all open-source LLMs, whereas the second comprises only the top-3 LLMs based on the training set. Besides, the best result of open-source LLMs on the test set of each benchmark is gathered to form the model entry “Best LLM” in Table 1.
Superiority of JiSi. The proposed JiSi framework exhibits promising performance, ranking first on average among all closed-source LLMs, open-source LLMs, router methods, and multi-agent methods. The performance gain over the most powerful open-source LLM (DeepSeek-V3.2-Speciale) is substantially 6.56%. Even compared with the state-of-the-art closed-source LLM Gemini-3-Pro, a +1.15% gain is achieved. Moreover, the router-only variant of JiSi, which directly adopts $\boldsymbol{A}_{1}^{fm}$ without the adaptive aggregation formulated in Eq. 5, outperforms other router methods, indicating that the proposed query-response mixed router can yield more suitable aggregatees for subsequent aggregation. Compared with the router-only variant, JiSi (w/o Adaptive Aggregation) benefits +1.41% from dynamic aggregator selection, and the proposed adaptive aggregation further improves the performance by +1.06%. Besides, with the additional aggregation, JiSi overpasses the theoretical performance bound of routing, i.e. “Best LLM”, by +1.6%, revealing enormous potential.
Efficiency of JiSi. As shown in Table 2, we report the cost of the proposed JiSi and the leading proprietary LLMs. It is worth noting that although JiSi utilizes more API calls due to multi-agent aggregation, it can remarkably reduce the total cost while maintaining superior performance compared with leading proprietary LLMs. For instance, JiSi surpasses Gemini-3-Pro and Grok-4 by +1.15% and +7.89%, while saving 53.23% and 81.68% cost. The reasons lie in two aspects: 1) the APIs of open-source LLMs are dramatically cheaper than proprietary LLMs. 2) The proposed adaptive routing-aggregation switch can mitigate the workload of aggregation and even simplify aggregation to routing. The strong trade-off between efficiency and effectiveness demonstrates the feasibility and economy of JiSi in practical implementation.
Is Self-Consistency Good Enough? The answer is “No”. It is worth noting that the multi-agent baseline “Self Consistency” (SC) seems to be performant enough, almost reaching the performance of the best closed-source LLM with a gap within 0.5%. However, the strong results from SC are based on an “oracle” setting that assumes the method can determine to which benchmark a new test question belongs and apply the most scored LLMs evaluated on the training set of this benchmark. It should be clarified that such a setting is practically infeasible, as the type of a given question is generally unknown. In contrast, the proposed JiSi is question-type-agnostic, using similarity metrics to construct a support set for LLM routing, which is practical and extensible. Moreover, the four-rollout SC, which costs almost equally as JiSi, does not provide a significant boost (just about 0.1%) compared with “Best LLM”, demonstrating the limitation of homogeneous LLM scaling.
Compared with Voting. Majority voting is widely used to make multi-agent decisions. However, it is fragile to LLM responses, especially when most collaborative LLMs cannot correctly handle a question. Voting may be biased toward a majority of incorrect answers, leading to worse results when scaling, as shown in Table 1 (-2.95% from 3 “Best LLMs” to “10 LLMs”). The proposed JiSi explores the minority of correct answers via fine-grained rerouting, thereby ensuring scalability while delivering extraordinary performance.
4.3 Scaling Ability
To verify the scalability of JiSi, we initialize a base LLM bank with five LLMs: DeepSeek-V3-0324, DeepSeek-R1-0528, Intern-S1, Qwen3-235B-A22B-2507, and Qwen3-235B-A22B-Thinking-2507. Subsequently, we introduce five additional open-source LLMs following their chronological release sequence: Kimi-K2-0905, DeepSeek-V3.1-Terminus, GLM-4.6, DeepSeek-V3.2-Thinking, and DeepSeek-V3.2-Speciale. This setup allows us to simulate the real-world evolution of the LLM ecosystem, resulting in five experimental configurations with model pool sizes $\{5,6,7,8,9,10\}$. All other hyperparameters remain the same as Sec. 4.1. To provide a multi-faceted evaluation, we categorize nine benchmarks into three specialized domains: 1) Reasoning & Knowledge: AIME, GPQA, MMLU-Pro, HLE, LiveMathBench. 2) Coding & Engineering: LiveCodeBench, SWE-bench. 3) General Chat & Factuality: Arena-Hard, SimpleQA. As illustrated in Fig. 5, JiSi demonstrates a robust and consistent performance gain across all three domains as the number of candidate LLMs increases. The consistent improvement across diverse benchmarks validates that JiSi is not merely a static ensemble but a dynamic system that scales with the progress of the open-source community. This empirical evidence supports a paradigm shift in intelligence enhancement, transitioning from the neural scaling law of parameters and training data towards a novel scaling law of multi-agent collaboration.
4.4 Distribution of LLM Selection
To further investigate the underlying routing logic of JiSi, we visualize the usage distribution of selected models across nine benchmarks in Fig. 6. The empirical distribution reveals three key insights.
Preference for High-Capacity LLMs.As illustrated in the overall distribution, JiSi performs a distinct preference for high-capacity models such as DeepSeek-V3.2-Speciale and Qwen3-235B-A22B-2507. By mainly routing queries to these top-performing open-source models with the highest individual average scores shown in Table 1, the system establishes a superior intelligence floor. This strategic selection ensures that the subsequent aggregation process operates on high-fidelity information, providing the necessary reasoning density for final high aggregation performance.
Task-Aware Specialization and Domain Expertise. JiSi exhibits remarkable sensitivity to the inherent requirements of different domains, effectively matching tasks with specialized experts. For instance, in reasoning-heavy tasks such as AIME and HLE, the router converged almost exclusively on DeepSeek-V3.2-Speciale. Meanwhile, for SimpleQA, which demands vast world knowledge and factuality, the system dynamically shifted its preference towards Qwen3-235B-A22B, which is the model with the fewest hallucinations and highest factual accuracy. It demonstrates that JiSi does not merely select the best overall model, but identifies the optimal matching between given query semantics and model expertise.
Abundant Usage Diversity against System Collapse. Crucially, JiSi avoids the pitfall of system collapse, where the system redundantly relies on a single LLM and degrades into a single model. This is most evident in Arena-Hard and MMLU-Pro, where the selection entropy is high, involving a diverse ensemble including DeepSeek-R1-0528, Kimi-K2, and Intern-S1. This heterogeneity is vital, which guarantees that the aggregator receives complementary perspectives, forming a prerequisite for the emergent performance gains. The preservation of this diverse candidate pool mitigates single-point biases and enhances the system’s generalization across multifaceted user prompts.
5 Conclusion
In this work, we propose JiSi, a training-free and scalable multi-agent system framework that leverages a minimalist “Route-and-Aggregate” topology. By incorporating three key innovations, including query-response mixed routing, support-set-based aggregator selection, and adaptive routing-aggregation switch, JiSi effectively orchestrates the collaboration of heterogeneous open-source LLMs while mitigating information noise. Experimental results across nine benchmarks demonstrate that JiSi not only significantly outperforms existing collective intelligence baselines but also surpasses the state-of-the-art proprietary model, Gemini-3-Pro, with an average gain of +1.15 while saving 53.23% cost. These findings validate that the strategic collaboration of specialized models provides a highly effective and scalable alternative to the continuous expansion of monolithic architectures. Our work paves the way for a potential and efficient trajectory towards artificial general intelligence.